what number should be added to -5/8 to get 5/9?????
 
 
 
 

Grade:8

5 Answers

Umakant Biswal
6 years ago
dear samiksha 
they are quite simple . 
just substract . 
5/9 – (-5/8) 
= 5/9 + 5/8 = 85/72 . 
HOPE IT CLEARS 
you can do that by taking x also 
kindly approve 
cheers 
Prashant
6 years ago
Suppose that number is x.
Then as per problem statement
x+(-5/8)= 5/9
x= 5/9+5/8
x=5(1/9+1/8)
x=5(17/72)=85/72=1.180555
C.Vishaal
6 years ago
Let the number to be added be x.
 
Then  -5/8 + x = 5/9
x = 5/9 + 5/8
x = 40 + 45/72
x = 85/72.
 
Therefore, the number to be added to – 5/8 in order to get +5/9 is 85/72.
 
Hope this answer helps.
praveen kumar
6 years ago
-5/8+x=5/9
x=5/9+5/8
x=(45+40)/72
x=85/72    the l.c.m of 9 and 8 is 72 so we get the denominator as 72
these are simple but clarify our doubts it’s good
Amit
6 years ago
-5/8+x=5/9x=5/9+5/8x=(45+40)/72x=85/72.